OpenStudy (anonymous):

Write a formula for the nth term of the sequence 7, 10, 13, 16

OpenStudy (tkhunny):

There are infinitely many such formulas. What sort would you like? IF the question had said "arithmetic sequence", THEN there would be a unique answer. The question didn't say that. Literally, write ANY nth term you can justify. n = 1 Term = 7 n = 2 Term = 10 n = 3 Term = 13 n = 4 Term = 16 Someone might point out that there seems to be a common difference of three so it MUST BE an arithmetic sequence with common difference of 3. You are free to assume this, but it is NOT the only solution. 10-7 = 3 13-10 = 3 16-13 = 3 n = 1 Term = 7 n = 2 Term = 10 = 7 + 3 n = 3 Term = 13 = 10 + 3 n = 4 Term = 16 = 13 + 3 n = 1 Term = 7 n = 2 Term = 10 = 7 + 3 n = 3 Term = 13 = 7 + 2*3 n = 4 Term = 16 = 7 + 3*3 I'm seeing one possible pattern. There are MANY others.

OpenStudy (anonymous):

Yes I know and this is an arithmetic series, the distance is 3, and so on...

OpenStudy (anonymous):

I'm not sure though, but would the answer be like an= 3n +4????

OpenStudy (tkhunny):

Who said it was an Arithmetic Sequence? Did you fail to report this as part of the problem statement? It is a VERY important piece of information.

OpenStudy (anonymous):

How is it not Arithmetic?

OpenStudy (anonymous):

ok nvm, but what's the answer, I get this stuff, I just need an answer

OpenStudy (anonymous):

ok well is this the answer, yes or no? an= 3n +4?

OpenStudy (anonymous):

tkhunny you the best

OpenStudy (anonymous):

guys, please i'm about to cry, just is this the answer yes or no? an= 3n +4 or would this be the answer? an = 7 +3 (n-1)? yes or no which one is it? or none of the above????

OpenStudy (tkhunny):

You do not get this stuff if you have argued what you just said. Try this function: \(f(x) = (1/6)(x^{3} - 6x^{2} + 29x + 18)\) Evaluate it for x = 1, x = 2, x = 3, and x = 4. Use synthetic division if you like. Then you tell me, "How can it not be arithmetic."

OpenStudy (anonymous):

hey I gtg, just please, just tell me the answer, if not you then somebody, please all i want is the answer PLEASE

OpenStudy (tkhunny):

You already created two correct answers. I gave you a third.